40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Nov 7 15:51:44 2012, Jenne, Update, Alignment, Jamie's tip tilt proposal 
    Reply  Wed Nov 7 17:20:01 2012, jamie, Update, Alignment, Jamie's tip tilt proposal 
       Reply  Wed Nov 7 23:22:45 2012, rana, Update, Alignment, Jamie's tip tilt proposal 
Message ID: 7686     Entry time: Wed Nov 7 23:22:45 2012     In reply to: 7684
Author: rana 
Type: Update 
Category: Alignment 
Subject: Jamie's tip tilt proposal 

 With Picos, we lose the ability to dither the input beam as well as align the beam with the IFO locked. And the active TT will still have hysteresis, but also actuators. Once in vacuum, I'm not sure how we adjust them - what's the error signal for PR2/PR3 ?

If the interferometer is aligned, why not just pump down now? I'm not sure that we have evidence of TT hysteresis issues once people stop touching them.

ELOG V3.1.3-